# On Arratia’s coupling and the Dirichlet law for the factors of a random integer

### v1

@@ 6. When $J_{x}$ and $J_{x}^{*}$ are different
$\mathbb{P}[J_{x}\neq J_{x}^{*}]\leqslant\mathbb{P}[\mathcal{E}_{1}^{c}]+\mathbb{P}[\mathcal{E}_{1}\cap\mathcal{E}_{2}^{c}]+\mathbb{P}[\mathcal{E}_{1}\cap\mathcal{E}_{3}^{c}].$
@@
Assume now further that $\mathcal{E}_{2}$ and $\mathcal{E}_{3}$ also occur. We claim that this implies $J_{x}=J_{x}^{*}$.
Let $B_{+}$ be the set of integers $i\geqslant 1$ such that $Y_{i}>e^{-\gamma}$, and let $B_{-}$ be the set of integers $i\leqslant-1$ such that $Y_{i}>e^{-\gamma}$. By our assumption that $\mathcal{E}_{2}\cap\mathcal{E}_{3}$ occurs and by the definition of $\eta$, we have
@@
Now, let $(T_{j}^{*},Q_{j}^{*})$ with $j\in\mathbb{Z}$ be the indexing of points of $\mathscr{R}^{*}$ given in Section 5, that is to say we have $T_{j}^{*}<T_{j+1}^{*}$ for all $j$, and such that $\prod_{j=0}^{\infty}Q_{j}^{*}>x\geqslant\prod_{j=1}^{\infty}Q_{j}^{*}$. Let $j_{0}\in\mathbb{Z}$ be such that $Q_{j_{0}}^{*}=e^{h(Y_{0})}$ (which exists because $Y_{0}>e^{-\gamma}$). Using relations (6.3) and (6.4), we find that
$\prod_{j>j_{0}}Q_{j}^{*}=\prod_{i\in B_{+}}e^{h(Y_{i})}=\prod_{i\geqslant 1}e^{h(Y_{i})}=J_{x}.$
On the other hand, we have $\prod_{j\geqslant j_{0}}Q_{j}^{*}=e^{h(Y_{0})}J_{x}$. Hence, using (6.2), we find that $\prod_{j>j_{0}}Q_{j}^{*}<x<\prod_{j\geqslant j_{0}}Q_{j}^{*}$. In particular, $j_{0}=0$ and thus $J_{x}^{*}=\prod_{j>j_{0}}Q_{j}^{*}=J_{x}$, as claimed. ∎
@@
Note that if $W_{k}>(\log_{2}x)^{-10}$ and $|S_{k}-\log x|\leqslant 2r_{0}+\log_{2}x$, then $|I_{(\log_{2}x)^{-10}}|\geqslant\log x-\log_{2}x-2r_{0}$. So, using Proposition 6.1, we find that
$\mathbb{P}\Big{[}|S_{k}-\log x|\leqslant\log_{2}x+2,\ W_{k}>(\log_{2}x)^{-10}\Big{]}$ $\leqslant\mathbb{P}\big{[}I_{(\log_{2}x)^{-10}}>\log x-\log_{2}x-2r_{0}\big{]}$
@@
Let $m_{0}$ be the largest integer such that $2^{m-1}r_{0}\leqslant\log_{2}x$. Hence, if $2r_{0}<R_{x}\leqslant\log_{2}x$, then there exists a unique integer $m\in[2,m_{0}]$ such that $2^{m-1}r_{0}<R_{x}\leqslant 2^{m}r_{0}$, in which case the condition $|S_{k}-\log x|\leqslant R_{x}+2r_{0}$ implies that $|S_{k}-\log x|\leqslant 2^{m}r_{0}+2r_{0}\leqslant 2^{m+1}r_{0}$. We conclude that
@@
$\mathbb{P}[\mathcal{E}_{1}^{c}]\leqslant\mathop{\sum\sum}_{\begin{subarray}{c}k\in\{0,1\}\\ 2\leqslant m\leqslant m_{0}\end{subarray}}\frac{4^{2-m}}{r_{0}^{2}}\mathbb{E}\bigg{[}\sum_{i>j>k}r(Y_{i})r(Y_{j})\cdot\mathbbm{1}_{|S_{k}-\log x|\leqslant 2^{m+1}r_{0}}\cdot\mathbbm{1}_{W_{k}\leqslant(\log_{2}x)^{-10}}\bigg{]}+O\bigg{(}\frac{1}{\log x}\bigg{)},$
@@
(6.5) $E(z)\coloneqq\mathbb{E}\bigg{[}\sum_{k\in\{0,1\}}\sum_{i>j>k}r(Y_{i})r(Y_{j})\cdot\mathbbm{1}_{|S_{k}-\log x|\leqslant z}\cdot\mathbbm{1}_{W_{k}\leqslant(\log_{2}x)^{-10}}\Big{]}\ll\frac{z}{\log x}$
@@
For the rest of the proof, we fix $z\in[0,4\log_{2}x]$. Given $t^{\prime\prime}>t^{\prime}>t$, let
@@
$E(z)\leqslant 2\cdot\mathbb{E}\bigg{[}\mathop{\sum\sum\sum}_{\begin{subarray}{c}(W,Y),(W^{\prime},Y^{\prime}),(W^{\prime\prime},Y^{\prime\prime})\in\mathscr{R}\\ W^{\prime\prime}>W^{\prime}>W\end{subarray}}r(Y^{\prime})r(Y^{\prime\prime})\cdot\mathbbm{1}_{|Y+Y^{\prime}+Y^{\prime\prime}+I_{W,W^{\prime},W^{\prime\prime}}-\log x|\leqslant z}\cdot\mathbbm{1}_{W\leqslant(\log_{2}x)^{-10}}\bigg{]}.$
Hence, using the Mecke equation (Proposition B.6), we find that
$E(z)\leqslant 2\mathop{\int\cdots\int}\limits_{\begin{subarray}{c}0<w<w^{\prime}<w^{\prime\prime}\\ w\leqslant(\log_{2}x)^{-10}\\ y,y^{\prime},y^{\prime\prime}\geqslant 0\\ y+y^{\prime}+y^{\prime\prime}\leqslant\log x+z\end{subarray}}r(y^{\prime})r(y^{\prime\prime})\mathbb{P}\Big{[}\big{|}I_{w}+y+y^{\prime}+y^{\prime\prime}-\log x\big{|}\leqslant z\Big{]}\frac{\,\mathrm{d}w\cdots\,\mathrm{d}y^{\prime\prime}}{e^{wy+w^{\prime}y^{\prime}+w^{\prime\prime}y^{\prime\prime}}},$
@@
$\mathbb{P}\Big{[}\big{|}I_{w}+y+y^{\prime}+y^{\prime\prime}-\log x\big{|}\leqslant z\Big{]}\leqslant e^{-w(\log x-y-y^{\prime}-y^{\prime\prime})}(e^{wz}-e^{-wz})\ll wze^{-w(\log x-y-y^{\prime})},$
since $w\leqslant(\log_{2}x)^{-10}$ and $z\leqslant 4\log_{2}x$ here. Consequently,
$\begin{split}E(z)&\ll\mathop{\int\cdots\int}\limits_{\begin{subarray}{c}0<w<w^{\prime}<w^{\prime\prime}\\ y,y^{\prime},y^{\prime\prime}\geqslant 0\\ y+y^{\prime}+y^{\prime\prime}\leqslant\log x+z\end{subarray}}r(y^{\prime})r(y^{\prime\prime})zwe^{-w\log x-(w^{\prime}-w)y^{\prime}-(w^{\prime\prime}-w^{\prime})(y^{\prime}+y^{\prime\prime})}\,\mathrm{d}w\cdots\,\mathrm{d}y^{\prime\prime}\\ &=\frac{z}{(\log x)^{2}}\iiiint\limits_{\begin{subarray}{c}t,y,y^{\prime},y^{\prime\prime}\geqslant 0\\ y+y^{\prime}+y^{\prime\prime}\leqslant\log x+z\end{subarray}}\frac{r(y^{\prime})r(y^{\prime\prime})}{y^{\prime}(y^{\prime}+y^{\prime\prime})}te^{-t}\,\mathrm{d}t\,\mathrm{d}y\,\mathrm{d}y^{\prime}\,\mathrm{d}y^{\prime\prime},\end{split}$
@@
$\mathbb{P}[\mathcal{E}_{1}\cap\mathcal{E}_{2}^{c}]\leqslant\mathbb{E}\bigg{[}\mathop{\sum\sum}_{(W,Y),(W^{\prime},Y^{\prime})\in\mathscr{R}}\mathbbm{1}_{1<W^{\prime}/W\leqslant 1+\eta/\min\{Y,Y^{\prime}\}^{2}}\cdot\mathbbm{1}_{Y^{\prime},Y>e^{-\gamma}}\cdot\mathbbm{1}_{I_{W}\in(\log x-Y,\log x]}\bigg{]}.$
We use Mecke’s equation as in the proof of (6.5) to get that
@@
$\mathbb{P}[\mathcal{E}_{1}\cap\mathcal{E}_{2}^{c}]$ $\ll\iiint\limits_{\begin{subarray}{c}w>0,\ y,y^{\prime}>e^{-\gamma}\\ y,y^{\prime}\leqslant\log x\end{subarray}}\frac{we^{-w\log x}}{\min\{y,y^{\prime}\}^{2}}\,\mathrm{d}w\,\mathrm{d}y\,\mathrm{d}y^{\prime}+\iiint\limits_{\begin{subarray}{c}w>0,\ y,y^{\prime}>e^{-\gamma}\\ y^{\prime}\leqslant\log x<y^{\prime}\end{subarray}}\frac{we^{-w(y+y^{\prime})}}{(y^{\prime})^{2}}\,\mathrm{d}w\,\mathrm{d}y\,\mathrm{d}y^{\prime}$
$\leqslant\iint\limits_{e^{-\gamma}<y,y^{\prime}\leqslant\log x}\frac{1}{\min\{y,y^{\prime}\}^{2}(\log x)^{2}}\,\mathrm{d}y\,\mathrm{d}y^{\prime}+\iint\limits_{\begin{subarray}{c}y,y^{\prime}>e^{-\gamma}\\ y^{\prime}\leqslant\log x<y^{\prime}\end{subarray}}\frac{1}{(yy^{\prime})^{2}}\,\mathrm{d}y\,\mathrm{d}y^{\prime}$
@@
For $x\geqslant 2$, we have $\mathbb{P}[J_{x}\neq J_{x}^{*}]\ll 1/\log x$.
@@
If we set $L(a)=\sum_{p|a}\log p$, then we have
$\mathop{\sum\sum}_{p,j:\ pj\in A}\frac{\log p}{\log x}=\sum_{a\in A}\frac{L(a)}{\log x}=\#A-\sum_{a\in A}\frac{\log(x/a)}{\log x}-\sum_{a\in A}\frac{\log a-L(a)}{\log x}.$
Since $\log(x/a)\geqslant 0$ for $a\leqslant x$, and we also have $L(a)\leqslant\log a$ for all $a$, relation (7.4) holds uniformly over all choices of $A\subseteq\mathbb{Z}\cap[1,x]$ if, and only if,
(7.5) $\sum_{a\leqslant x}\log(x/a)\ll x\quad\text{and}\quad\sum_{a\leqslant x}(\log a-L(a))\ll x.$
The first estimate follows readily by partial summation (or by Stirling’s formula). The second one is a simple consequence of the identity $\log a-L(a)=\sum_{p^{\nu}\|a,\,\nu\geqslant 2}(\nu-1)\log p$. This establishes (7.5), and hence it completes the proof of Proposition 2.4. ∎
@@ 8. Theorem 2 for the probabilistic model
Let $\mathbf{V}=(V_{1},V_{2},\ldots)$ be a Poisson–Dirichlet process and let $(C_{i})_{i\geqslant 1}$ be a sequence of i.i.d. random variables (also independent of $\mathbf{V}$) with $\mathbb{P}[C_{i}=j]=\alpha_{j}$ for all $i\geqslant 1$. Then the random vector
@@
for $\text{Re}(s)<1$. This coincides with the moment generating function of the distribution $\text{Gamma}(\alpha_{m},1)$. We thus deduce that the vector

### v2

@@ 6. When $J_{x}$ and $J_{x}^{*}$ are different
$\mathbb{P}\big{[}J_{x}\neq J_{x}^{*}\big{]}\leqslant\mathbb{P}[\mathcal{E}_{1}^{c}]+\mathbb{P}[\mathcal{E}_{1}\cap\mathcal{E}_{2}^{c}]+\mathbb{P}[\mathcal{E}_{1}\cap\mathcal{E}_{3}^{c}].$
@@
Assume now further that $\mathcal{E}_{2}$ and $\mathcal{E}_{3}$ also occur. We claim that this implies $J_{x}=J_{x}^{*}$. Let $B_{+}$ be the set of integers $i\geqslant 1$ such that $Y_{i}>e^{-\gamma}$, and let $B_{-}$ be the set of integers $i\leqslant-1$ such that $Y_{i}>e^{-\gamma}$. By our assumption that $\mathcal{E}_{2}\cap\mathcal{E}_{3}$ occurs and by the definition of $\eta$, we have
@@
Now, let $K\geqslant 0$ and $(T_{j}^{*},Q_{j}^{*})$ with $j\in\mathbb{Z}_{\leqslant K}$ be the indexing of points of $\mathscr{R}^{*}$ given in Section 5, that is to say we have $T_{j-1}^{*}<T_{j}^{*}$ for all $j\in\mathbb{Z}_{\leqslant K}$, and such that $\prod_{j=0}^{K}Q_{j}^{*}>x\geqslant\prod_{j=1}^{K}Q_{j}^{*}$. Let $j_{0}\in\mathbb{Z}_{\leqslant K}$ be such that $T_{j_{0}}^{*}=W_{0}Y_{0}/h(Y_{0})$ (which exists because $Y_{0}>e^{-\gamma}$). Using relations (6.3) and (6.4), we find that
$\prod_{j_{0}<j\leqslant K}Q_{j}^{*}=\prod_{i\in B_{+}}e^{h(Y_{i})}=\prod_{i\geqslant 1}e^{h(Y_{i})}=J_{x}.$
On the other hand, we have $\prod_{j_{0}\leqslant j\leqslant K}Q_{j}^{*}=e^{h(Y_{0})}J_{x}$. Hence, using (6.2), we find that $\prod_{j_{0}<j\leqslant K}Q_{j}^{*}<x<\prod_{j_{0}\leqslant j\leqslant K}Q_{j}^{*}$. In particular, $j_{0}=0$ and thus $J_{x}^{*}=\prod_{j_{0}<j\leqslant K}Q_{j}^{*}=J_{x}$, as claimed. ∎
@@
Note that if there exists $k\in\mathbb{Z}$ with $W_{k}>(\log_{2}x)^{-10}$ and $|S_{k}-\log x|\leqslant\log_{2}x+2r_{0}$, then $I_{(\log_{2}x)^{-10}}\geqslant\log x-\log_{2}x-2r_{0}$. So, using Proposition 6.1, we find that
$\mathbb{P}\Big{[}\exists k:|S_{k}-\log x|\leqslant\log_{2}x+2r_{0},\ W_{k}>(\log_{2}x)^{-10}\Big{]}$ $\leqslant\mathbb{P}\big{[}I_{(\log_{2}x)^{-10}}>\log x-\log_{2}x-2r_{0}\big{]}$
@@
Let $m_{0}$ be the largest integer such that $2^{m_{0}-1}r_{0}\leqslant\log_{2}x$. Hence, if $2r_{0}<R_{x}\leqslant\log_{2}x$, then there exists a unique integer $m\in[2,m_{0}]$ such that $2^{m-1}r_{0}<R_{x}\leqslant 2^{m}r_{0}$, in which case the condition $|S_{k}-\log x|\leqslant R_{x}+2r_{0}$ implies that $|S_{k}-\log x|\leqslant 2^{m}r_{0}+2r_{0}\leqslant 2^{m+1}r_{0}$. We conclude that
@@
$\mathbb{P}[\mathcal{E}_{1}^{c}]\leqslant\mathop{\sum\sum}_{\begin{subarray}{c}k\in\{0,1\}\\ 2\leqslant m\leqslant m_{0}\end{subarray}}\frac{4^{2-m}}{r_{0}^{2}}\mathbb{E}\bigg{[}\mathop{\sum\sum}_{i,j:\ i>j>k}r(Y_{i})r(Y_{j})\cdot\mathbbm{1}_{|S_{k}-\log x|\leqslant 2^{m+1}r_{0}}\cdot\mathbbm{1}_{W_{k}\leqslant(\log_{2}x)^{-10}}\bigg{]}+O\bigg{(}\frac{1}{\log x}\bigg{)},$
@@
(6.5) $E(z)\coloneqq\mathbb{E}\bigg{[}\mathop{\sum\sum\sum}_{\begin{subarray}{c}i,j,k\in\mathbb{Z}:\ i>j>k\end{subarray}}r(Y_{i})r(Y_{j})\cdot\mathbbm{1}_{|S_{k}-\log x|\leqslant z}\cdot\mathbbm{1}_{W_{k}\leqslant(\log_{2}x)^{-10}}\Big{]}\ll\frac{z}{\log x}$
@@
For the rest of the proof, we fix $z\in[0,4\log_{2}x]$. Given $t^{\prime\prime}>t^{\prime}>t>0$, let
@@
$E(z)\leqslant\mathbb{E}\bigg{[}\mathop{\sum\sum\sum}_{\begin{subarray}{c}(W,Y),(W^{\prime},Y^{\prime}),(W^{\prime\prime},Y^{\prime\prime})\in\mathscr{R}\\ W^{\prime\prime}>W^{\prime}>W\end{subarray}}r(Y^{\prime})r(Y^{\prime\prime})\cdot\mathbbm{1}_{|Y+Y^{\prime}+Y^{\prime\prime}+I_{W,W^{\prime},W^{\prime\prime}}-\log x|\leqslant z}\cdot\mathbbm{1}_{W\leqslant(\log_{2}x)^{-10}}\bigg{]}.$
Hence, using the Mecke equation (Proposition B.6) and the fact that $I_{t,t^{\prime},t^{\prime\prime}}$ has the same distribution as $I_{t}$, we find that
$E(z)\leqslant\mathop{\int\cdots\int}\limits_{\begin{subarray}{c}0<w<w^{\prime}<w^{\prime\prime}\\ w\leqslant(\log_{2}x)^{-10}\\ y,y^{\prime},y^{\prime\prime}\geqslant 0\\ y+y^{\prime}+y^{\prime\prime}\leqslant\log x+z\end{subarray}}r(y^{\prime})r(y^{\prime\prime})\mathbb{P}\Big{[}\big{|}I_{w}+y+y^{\prime}+y^{\prime\prime}-\log x\big{|}\leqslant z\Big{]}\frac{\,\mathrm{d}w\cdots\,\mathrm{d}y^{\prime\prime}}{e^{wy+w^{\prime}y^{\prime}+w^{\prime\prime}y^{\prime\prime}}},$
@@
$\mathbb{P}\Big{[}\big{|}I_{w}+y+y^{\prime}+y^{\prime\prime}-\log x\big{|}\leqslant z\Big{]}\leqslant e^{-w(\log x-y-y^{\prime}-y^{\prime\prime})}(e^{wz}-e^{-wz})\ll wze^{-w(\log x-y-y^{\prime}-y^{\prime\prime})},$
since $w\leqslant(\log_{2}x)^{-10}$ and $z\leqslant 4\log_{2}x$. Consequently,
$\begin{split}E(z)&\ll\mathop{\int\cdots\int}\limits_{\begin{subarray}{c}0<w<w^{\prime}<w^{\prime\prime}\\ y,y^{\prime},y^{\prime\prime}\geqslant 0\\ y+y^{\prime}+y^{\prime\prime}\leqslant\log x+z\end{subarray}}r(y^{\prime})r(y^{\prime\prime})zwe^{-w\log x-(w^{\prime}-w)(y^{\prime}+y^{\prime\prime})-(w^{\prime\prime}-w^{\prime})y^{\prime\prime}}\,\mathrm{d}w\cdots\,\mathrm{d}y^{\prime\prime}\\ &=\frac{z}{(\log x)^{2}}\iiiint\limits_{\begin{subarray}{c}t,y,y^{\prime},y^{\prime\prime}\geqslant 0\\ y+y^{\prime}+y^{\prime\prime}\leqslant\log x+z\end{subarray}}\frac{r(y^{\prime})r(y^{\prime\prime})}{y^{\prime\prime}(y^{\prime}+y^{\prime\prime})}te^{-t}\,\mathrm{d}t\,\mathrm{d}y\,\mathrm{d}y^{\prime}\,\mathrm{d}y^{\prime\prime},\end{split}$
@@
$\mathbb{P}[\mathcal{E}_{1}\cap\mathcal{E}_{2}^{c}]\leqslant\mathbb{E}\bigg{[}\mathop{\sum\sum}_{(W,Y),(W^{\prime},Y^{\prime})\in\mathscr{R}}\mathbbm{1}_{1<W^{\prime}/W\leqslant 1+\eta/\min\{Y,Y^{\prime}\}^{2}}\cdot\mathbbm{1}_{Y^{\prime},Y>e^{-\gamma}}\cdot\mathbbm{1}_{I_{W,W^{\prime}}+Y^{\prime}\in(\log x-Y,\log x]}\bigg{]}$
with $I_{t,t^{\prime}}\coloneqq\sum_{(W,Y)\in\mathscr{R},\ W\in\mathbb{R}_{>t}\setminus\{t^{\prime}\}}Y$. We use Mecke’s equation as in the proof of (6.5) to get that
@@
$\mathbb{P}[\mathcal{E}_{1}\cap\mathcal{E}_{2}^{c}]$ $\ll\iiint\limits_{\begin{subarray}{c}w>0,\ y,y^{\prime}>e^{-\gamma}\\ y,y^{\prime}\leqslant\log x\end{subarray}}\frac{we^{-w\log x}}{\min\{y,y^{\prime}\}^{2}}\,\mathrm{d}w\,\mathrm{d}y\,\mathrm{d}y^{\prime}+\iiint\limits_{\begin{subarray}{c}w>0,\ y,y^{\prime}>e^{-\gamma}\\ y^{\prime}\leqslant\log x<y\end{subarray}}\frac{we^{-w(y+y^{\prime})}}{(y^{\prime})^{2}}\,\mathrm{d}w\,\mathrm{d}y\,\mathrm{d}y^{\prime}$
$\leqslant\iint\limits_{e^{-\gamma}<y,y^{\prime}\leqslant\log x}\frac{1}{\min\{y,y^{\prime}\}^{2}(\log x)^{2}}\,\mathrm{d}y\,\mathrm{d}y^{\prime}+\iint\limits_{\begin{subarray}{c}y,y^{\prime}>e^{-\gamma}\\ y^{\prime}\leqslant\log x<y\end{subarray}}\frac{1}{(yy^{\prime})^{2}}\,\mathrm{d}y\,\mathrm{d}y^{\prime}$
@@
For $x\geqslant 2$, we have $\mathbb{P}\big{[}J_{x}\neq J_{x}^{*}\big{]}\ll 1/\log x$.
@@
If we set $L(a)\coloneqq\sum_{p|a}\log p$, then we have
(7.5) $\mathop{\sum\sum}_{p,j:\ pj\in A}\frac{\log p}{\log x}=\sum_{a\in A}\frac{L(a)}{\log x}=\#A-\sum_{a\in A}\frac{\log(x/a)+(\log a-L(a))}{\log x}$
The quantity $\log(x/a)$ is non-negative whenever $a\leqslant x$. Furthermore, the integer $a/(\prod_{p|a}p)$ is always a divisor of $s(a)$, thus $0\leqslant\log a-L(a)\leqslant\log s(a)$. Therefore,
$\sum_{a\in A}\frac{\log(x/a)+(\log a-L(a))}{\log x}$ $\ll\sum_{a\leqslant x}\frac{\log(x/a)+\log s(a)}{\log x}$
$=\frac{\left\lfloor x\right\rfloor}{\log x}\cdot\mathbb{E}\big{[}\log(x/N_{x})+\log s(N_{x})\big{]}$
$\ll\frac{x}{\log x}$
by Lemma 2.2. With this inequality combined with (7.5), we establish (7.4), and hence it completes the proof of Proposition 2.4. ∎
@@ 8. Theorem 2 for the probabilistic model
Let $\mathbf{V}=(V_{1},V_{2},\ldots)$ be a Poisson–Dirichlet process, let $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{k})\in\Delta^{k-1}$ such that $\alpha_{i}>0$ for all $1\leqslant i\leqslant k$, and let $(C_{i})_{i\geqslant 1}$ be a sequence of i.i.d. random variables (also independent of $\mathbf{V}$) with $\mathbb{P}[C_{i}=j]=\alpha_{j}$ for all $i\geqslant 1$. Then the random vector
@@
for $\text{Re}(s)<1$. This coincides with the moment generating function of $\text{Gamma}(\alpha_{m},1)$ distribution. We thus deduce that the vector